OpenStudy (anonymous):

Find the equation of the line tangent to the graph of y=4e^x at x=3. tangent line y=

OpenStudy (shubhamsrg):

any attempt ?

OpenStudy (anonymous):

well the tangent line is the derivative right so i got just the same thing 4e^x im not sure if i am right

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

now evaluate it at x=3 so you get 4e^3

OpenStudy (shubhamsrg):

thats the slope for any x. just substitute x=3 and you'll have the slope of the curve at x=3 for the equation, you already know the slope, you'd just need a point . You already know that point, don;t you ?

OpenStudy (anonymous):

well i plugged in that for my answer and i got it incorrect so i was a bit confuse

OpenStudy (anonymous):

Can you find the derivative of y=4(e^x)?

OpenStudy (anonymous):

Also, are you looking for point slope form, or slope intercept form?

OpenStudy (anonymous):

well i thought that the derivative was just the 4e^x but im not sure and it just tells me to find the tangent equation

OpenStudy (campbell_st):

well by 1st task in this question is to for the y value of the point where x = 3 so \[y = 4e^3\] so the point is \[(3, 4e^3)\] next find the 1st derivative, which you have done \[f'(x) = 4e^x\] find the slope when x = 3 \[m = 4e^3\] so using the equation of the line y = mx + b substitute and you get \[4e^3 = 4e^3(3) + b\] all you need to do is solve for b, the y- intercept \[4e^3 = 12e^3 + b\] then your equation will be \[y = 4e^3x + b\] hope this helps.

OpenStudy (anonymous):

so i got 4e^3x+4 is that correct

OpenStudy (e.mccormick):

Why +4?

OpenStudy (e.mccormick):

\(4e^3 = 12e^3 + b \implies 4e^3-12e^3 = b\)

OpenStudy (anonymous):

ok first calculate derivative 0f Y that is = 4e^x that is the slope of tangent at x=3 slpoe =4e^3 tangent is passing thru x=3 so find y point y=4e^x at x=3 y = 4e^3 (3,4e^3) so the eqation of line Y=mX+c put x, y ,slope in it 4e^3=4e^3*3+c c=-8e^3 y=4e^3x-8e^3